OpenStudy (anonymous):

For the curve y=2-12x+9x^2-2x^3: i) Find the intervals of increase and decrease. ii) Find the local maximum and minimum values. iii) Find the intervals of concavity. iv) Find any points of inflection. v) Sketch the curve.

OpenStudy (accessdenied):

The first step would just be finding the derivative, since most of the questions can be answered with information about it.

OpenStudy (anonymous):

y'=-6x^2+18x+12

OpenStudy (cwrw238):

y' = -12 + 18x - 6x^2

OpenStudy (anonymous):

**-12

OpenStudy (cwrw238):

solve y' = 0 to find max/minm

OpenStudy (anonymous):

how to do you solve y'=0?

OpenStudy (accessdenied):

Set "-12 + 18x - 6x^2 = 0" solve for x

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I really need to understand how to do part i) if you guys can help me out with that one first

OpenStudy (accessdenied):

well, finding the intervals also involves finding the 0's of the derivative. We'll use the points where the derivative is 0 as reference points and test x-values around them to find if the function is increasing or decreasing at those points (since when y' = 0, the graph could either change signs or stay the same, so these points are important to look at)

OpenStudy (anonymous):

ok to do y'=o what exactly do I do

OpenStudy (accessdenied):

We'll write out what we have for y' -6x^2 + 18x - 12 = 0 (written in proper general form) we can factor out a monomial -6 and divide it off since its just a constant x^2 - 3x + 2 = 0 then we factor the new expression since -1 * -2 = 2 and -1 + -2 = -3 (x - 1)(x - 2) = 0 ab = 0; a = 0 or b = 0 x - 1 = 0 and x - 2= 0 x = 1, x = 2

OpenStudy (anonymous):

Ok I get that

OpenStudy (anonymous):

so it is (x-1)(x+2)? right

OpenStudy (anonymous):

good Question. :)

OpenStudy (accessdenied):

it factors to (x-1)(x-2) = 0 we solve for the x-values that will get us 0 x - 1 = 0 x = 1 x - 2 = 0 x = 2 These values will be important to mark out our intervals.  We'd test x-values around those intervals in the derivative to determine if it is postivie or negative for that interval. ie. we could use x=0, x=1/2, and x=3

OpenStudy (accessdenied):

* that would tell us if that interval is increasing or decreasing, since a positive derivative indicates increase and a negative derivative indicates decrease.

OpenStudy (anonymous):

ok so where do we substitute these test values and how do we get the test values?

OpenStudy (accessdenied):

We substitute them into the derivative y' = -6x^2 + 18x - 12 and find the y' value They're pretty much our own choice. We can choose any values as long as they're not x=1 or x=2 (we already know the derivative is 0 there, no need to test them.) If you wanted, we could use x=26 rather than x=3 but it wouldn't be as nice to calculate. :)

OpenStudy (anonymous):

when do we know when to stop testing?

OpenStudy (accessdenied):

we only need to test one value of x in each interval x < 1, 1 < x < 2, and 2 < x

OpenStudy (anonymous):

oo so because its a parabola the intervals would be positive for 1<x<2 and negative for x<1 and x>2 right?

OpenStudy (anonymous):

My bad its -2

OpenStudy (accessdenied):

yeah, that would be correct graphically, you can tell pretty easily my method is just the analytical way, if you didn't have he graph. it is actually positive 2, not -2. since (x-1)(x-2) given x=2 => (2-1)(2-2) = 1*0 = 0 works

OpenStudy (anonymous):

i thought it was (x-1)(x+2)?

OpenStudy (anonymous):

oo ye nvm ur right

OpenStudy (accessdenied):

kk so, we have the intervals of change. x<1 decreasing 1<x<2 increasing x>2 decreasing To identify the local maxima and minima, we find them by using those points we found for the intervals (x=1 and x=2) and look at the intervals just around the points if the function goes from decreasing to increasing, we have a minimum value at the point if the function goes from increasing to decreasing, we have a maximum value at the poin

OpenStudy (anonymous):

but since its a parabola wont it only have 1 min/max?